Atlas tile calorimeter wbs 3 4
Download
1 / 21

ATLAS Tile Calorimeter WBS 3.4 - PowerPoint PPT Presentation


  • 109 Views
  • Uploaded on

ATLAS Tile Calorimeter WBS 3.4. Larry Price ANL. M&O Status and Plans Toronto August 3, 2005. Construction Project (Installation). Installation Schedule in Pit. Barrel March-December 2004 EBC: August-November 2005 (Oct-Nov) EBA: December-March 2006 (Dec-Jan).

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' ATLAS Tile Calorimeter WBS 3.4' - michon


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
Atlas tile calorimeter wbs 3 4

ATLAS Tile CalorimeterWBS 3.4

Larry Price

ANL

M&O Status and Plans

TorontoAugust 3, 2005


Construction project installation
Construction Project (Installation)


Installation schedule in pit
Installation Schedule in Pit

  • Barrel March-December 2004

  • EBC: August-November 2005 (Oct-Nov)

  • EBA: December-March 2006 (Dec-Jan)


Assembly/commissioning schedule (Version 7, still official)

  • Comments:

  • -Probable 4 weeks delay on LB movement (not official)=> delay EB assembly+LB commissioning

  • EC toroid A that follows EBA assembly is not available in time.

  • When EB is pushed inside toroid for ID installation or solenoid magnetic field test, the dist LB-EB~ 4m and minivan-scaffolding still in place =>can access EB/B fingers (to be looked in more detail).


Schedule for barrel Commissioning phase 1 active Tile WPs ( based on V7 schedule) 20 June05

This table does not taken into account that unofficially (most probably) the start of commissioning by sectors planned for 7 Sept-15Nov05 will be delayed by

3-4 weeks due to delay of the toroid installation/barrel movement




Tilecal front end electronics repairs on hf truck sept oct 04

TileCal Front-End Electronics Repairs on HF Truck Sept-Oct 04

-Tests on: TTC, CANbus, HV, laser fibres via LED, BCID @ 100 kHz L1A, readout noise, …

- Replaced failing components

- Will continue in Jan-Feb when OPAL cradle will be removed


Wp 4 1 in situ fe electronics certification with mobidick status in june05 claudio fabrice
WP 4.1: In situ FE electronics certification with Mobidick 04 (status in June05; Claudio,Fabrice)

  • Remaining tasks for z=0 (before starting commissioning, no access now):

  • 14 Super Drawers to be repaired ( 7 out of 64 modules)

  • 2 SD to be inserted

  • Repair/certify FE in cases of failures are found with MobiDAQ

  • Finish this week last 4 modules certification (without fingers)

  • Next step in July+Aug (Nikos): perform FE stability tests in all modules using MobiDAQ +temporary cables (~ 8 SD running ~ 3 days with portable cooling).


WP4.2: FE tests with MOBIDAQ 04(Gerolf,Iacopo,Amir,Oleg,Giorgi,Andrea)

(8 drawers at the time from USA15+temporarly cables/fibres)

MobiDAQ

  • MOBIDAQ in pit for many weeks and working in stable mode for a few days.

  • Still need to include LED test to send light to PMTs (will replace laser system for fibres QC)

Cosmic receiver setup

PC (MobiDAQ

interface)


Cosmic rays expert week 20-28 June (Richard et al.) 04

8 superdrawers (2 top +2 bottom modules)

8 sets of temporary cables (TTC/GLINK, CANbus, trigger, laser/LED)

8 temporary LVPS

2 temporary cooling units

MobiDAQ in USA15

2 Coincidence boards (Chicago)

Details:http://agenda.cern.ch/askArchive.php?base=agenda&categ=a053788&id=a053788s1t6%2Ftransparencies%2FTEUSCHER_Cosmic_Ray_Expert_Week_Summary_June_2005.ppt



Next expert weeks integration runs
Next expert weeks/integration runs: 04

  • July ’05: ROD Expert week in USA15

    • Read out 8 superdrawers with final ROD Motherboard

    • Debug DAQ/ROD

  • Dec ’05/Jan ‘06: Begin cosmic-ray running on barrel

    • Following installation of final services and Low Voltage PS at z=0

    • Up to 12 top + 12 bottom modules

  • 2006: Continue periodic runs until global ATLAS cosmic run in March ‘07

    • Integrate with LAr barrel after cool-down (~April ‘06)

    • Repeat with TileCal EBA / EBC / HEC / FCAL when ready in ‘06



Mbts counters
MBTS counters 04

2 segments in h; inner segment covers pixel detectors


Principal us m o work in 2005
Principal US M&O 04Work in 2005

  • Validation, calibration, and repair of modules and electronics

    • Adjustments and modifications to accommodate adjacent systems, e.g., LAr and Cryostat

    • Establishment of test/repair facility

  • Work with TC on services validation

  • Cosmic Ray Commissioning

    • Cosmic ray trigger system

    • “First ATLAS Data from the underground area”

  • Performance monitoring and calibration software development

  • Setup for cosmic ray commissioning and performance measurement

  • Vapor shields for central and forward cryostats

  • MBTS Scintillators

  • Validation and testing of LVPS

  • Initial integration with TDAQ and extensive exercise of interfaces (MobiDAQ)

  • Establishment of conditions and other databases

  • Design and prototyping of MBTS Counters


Plans for us m o work in 2006
Plans for US M&O 04Work in 2006

  • Complete MBTS Counters

  • Barrel integration with TDAQ + DCS and extensive exercise of interfaces

  • Barrel commissioning

    • Locally by sectors

    • Cosmics with TDAQ readout

  • EB Commissioning (as above)

  • Final mechanical and light integrity validation of EBA modules before installation in pit

  • Software development

    • Complete initial calibration, diagnostic, and status reporting software

    • Database development

  • Establish and exercise the procedure for regular diagnostics and calibrations

  • Validation and testing of LVPS

  • Identify and carry out needed electronics repairs

    • Completion of test/repair facility

  • Understand calorimeter performance and integrate with LAr and physics studies



Issues
Issues 04

  • Completion of LVPS

  • Congestion in the pit

    • Tile Extended Barrel assembly

    • Muon Barrel assembly

    • Tile Barrel commissioning

  • Will need probably to work in 2 shifts/day in Sept->Dec with several teams working in parallel.

  • Consequently need:

    • Big teams of physicists/tech/eng. at CERN during second half 2005 and first half of 2006.

    • Commission each Tile sub-system in standalone as much as possible before end August in USA15/UX15 with temporary cables

    • Good coordination within Tile sub-systems and with other ATLAS systems (through TC).

TC


ad